OpenStudy (anonymous):

Which fraction is the same as 3.4

OpenStudy (anonymous):

..... put it into the calculator 17/5

OpenStudy (anonymous):

3.4=3+0.4=3+4/10=3+2/5=15/5+2/5=17/5

OpenStudy (mathstudent55):

Here is another way of doing it. The 4 is in the tenths place, so this number can be read as 34 tenths, or 34/10. Then just reduce the fraction by dividing the numerator and denominator by 2 to get 17/5

OpenStudy (anonymous):

that's a good way to solve it.